### v1

# IndiSeek learns information-guided disentangled representations

Abstract
Learning disentangled representations is a fundamental task in multi-modal learning. In modern applications such as single-cell multi-omics, both shared and modality-specific features are critical for characterizing cell states and supporting downstream analyses. Ideally, modality-specific features should be independent of shared ones while also capturing all complementary information within each modality. This tradeoff is naturally expressed through information-theoretic criteria, but mutual-information–based objectives are difficult to estimate reliably, and their variational surrogates often underperform in practice. In this paper, we introduce IndiSeek, a novel disentangled representation learning approach that addresses this challenge by combining an independence-enforcing objective with a computationally efficient reconstruction loss that bounds conditional mutual information. This formulation explicitly balances independence and completeness, enabling principled extraction of modality-specific features. We demonstrate the effectiveness of IndiSeekon synthetic simulations, a CITE-seq dataset and multiple real-world multi-modal benchmarks.
1 Introduction
The growing availability of multi-modal data has broadened the scope of representation learning. One notable example is the development of single-cell multi-omics technologies in genomics (Teichmann and Efremova,, 2020). For instance, CITE-seq (Cellular Indexing of Transcriptomes and Epitopes by Sequencing) (Stoeckius et al.,, 2017) enables simultaneous measurement of gene expression via single-cell RNA sequencing and protein abundance via antibody-derived tags (ADTs). Integrative analysis of both modalities at the single-cell resolution has improved tasks such as cell state delineation and enabled cell atlas construction and querying at finer granularity. Multi-modal data is also abundant at the intersection of computer vision, natural language processing, and audio processing, where combining information from different sources for representation learning forms the backbone of many groundbreaking progresses (Radford et al.,, 2021; Baltrušaitis et al.,, 2018; Jia et al.,, 2021; Akbari et al.,, 2021; Liang et al.,, 2024).
The central question of multi-modal learning is how to extract informative and interpretable representations from complex multi-modal data. This involves two complementary goals:
  1. (1)
    Extracting interdependence (or shared information) across modalities;
  2. (2)
    Extracting modality-specific information unique to each modality.
Shared features enhance predictive power across modalities by capturing cross-modal dependence. From an information-theoretic perspective, they should contain all information common across modalities. zhe pu At the same time, preserving modality-specific diversity is equally important. In single-cell multi-omics, modality-specific features often correspond to unique biological signals that aid cell type identification and the discovery of novel or rare cell populations (Hao et al.,, 2021; Caron et al.,, 2025). After all, if all the information in a certain modality can be inferred from other modalities, there is little reason to measure it in the first place. Finally, disentangling shared and modality-specific components improves interpretability: shared features summarize cross-modal information, while modality-specific features capture complementary signals.
Methodologically, shared information is usually extracted by maximizing cross-modal mutual information (Tosh et al.,, 2021; Sridharan and Kakade,, 2008), with contrastive methods like CLIP (Radford et al.,, 2021) widely adopted in practice. Under certain conditions, CLIP-learned representations have been shown to be not only sufficient (i.e., capturing all shared information), but also minimally sufficient (i.e., containing minimal extra information) (Gui et al.,, 2025; Oko et al.,, 2025; Lin and Mei,, 2025) under certain conditions.
In contrast, identifying modality-specific features, also known as disentangled representation learning, has become an active research area (Fischer,, 2020; Liang et al.,, 2023; Liu et al.,, 2023; Wang et al., 2024a, ; Dufumier et al.,, 2024; Wang et al., 2024b, ). While shared features should be minimally sufficient, modality-specific features must both remain independent of (i.e., disentangled from) shared features and capture all complementary information within each modality (Liang et al.,, 2023; Wang et al., 2024a, ; Wang et al., 2024b, ). From an information-theoretic perspective, this requires (1) simultaneously minimizing the mutual information between shared and modality-specific features and (2) maximizing the mutual information between data and their union. Multiple state-of-the-art methods (Liang et al.,, 2023; Wang et al., 2024a, ) tackle this problem from such a perspective, but differ in how they approximate the underlying mutual information. Since such quantities are difficult to estimate with finite samples, these methods often trade off independence against completeness. As we show in the motivating examples below, this tradeoff limits their ability to achieve both goals simultaneously.
1.1 Motivating examples and limitations of SOTA
To demonstrate the limitation of the state-of-the-art (SOTA) methods in disentangled learning, we consider two simulated settings. To isolate the challenge of modality-specific feature extraction, we assume the shared features $C_{1}=f_{1}(X_{1})\in\mathbb{R}^{d_{c}}$ are provided by an oracle. The task is then to learn modality-specific features from a single modality $X_{1}\in\mathbb{R}^{d_{1}}$, with $d_{1}=6$ and $d_{c}=2$.
  • Setting 1: Let the observed data be iid copies of $X_{1}{\sim}{\mathcal{N}}(0,I_{d_{1}})$. For each $x\in\mathbb{R}^{d_{1}}$, define the shared representation map
    $f_{1}(x)=0.5A_{f}x+0.2\sin(A_{f}x)+0.2(A_{f}x)^{3}\quad\text{with}\quad A_{f}=(\mathbf{I}_{d_{c}},\mathbf{O})\in\mathbb{R}^{d_{c}\times d_{1}}.$ (1)
    Here, the sine and cubic functions are applied entrywise. The ideal modality-specific features are the last four coordinates of $X_{1}$, which contain all remaining information while being independent of $C_{1}$.
  • Setting 2: Let the observed data be iid copies of $X_{1}$ where $(X_{1})_{\{1,2,5,6\}}\sim{\mathcal{N}}(0,\mathbf{I}_{4})$, $(X_{1})_{3}=0.2\times((X_{1})_{1}+(X_{1})_{2})$, and $(X_{1})_{4}=(X_{1})_{1}\times(X_{1})_{2}$. Thus, the third and fourth coordinates are deterministic functions of the first two, while the others are independent. Let $C_{1}=(X_{1})_{1:d_{c}}$. Here, the ideal modality-specific features are the last two coordinates.
For both settings, given $C_{1}$, our goal is to learn a map $h_{1}:\mathbb{R}^{d_{1}}\to\mathbb{R}^{p_{1}}$ with $p_{1}=10$ such that $h_{1}(X_{1})$ gives the modality-specific features. All neural networks are five-layer ReLU MLPs with width $100$, trained on $10000$ samples. Ablation studies beyond Gaussian distributions are presented in the appendix.
For any learned map $h_{1}$, we quantify the importance of each coordinate of $X_{1}$ in determining its output via gradients based on feature masking [^1] See Appendix A.1 for details on this feature importance metric. on an independent test set of size $1000$, following standard model-free importance measures (Robnik-Šikonja and Kononenko,, 2008; Zeiler and Fergus,, 2014; Li et al.,, 2016)
(a) Factorized CL.
(b) InfoDisen.
(c) Our method (IndiSeek).
Figure 1: Importance of learned modality-specific features: Setting 1.
(a) Factorized CL.
(b) InfoDisen.
(c) Our method (IndiSeek).
Figure 2: Importance of learned modality-specific features: Setting 2.
We present the average coordinate importance metrics in learned modality-specific representation maps over $50$ simulation runs in Figures 1 and 2 for three different learning algorithms in the two settings, respectively. In both figures, the left panels report results of the SOTA approach FactorizedCL in Liang et al., (2023) (without self-supervision), the middle panels report results of the SOTA approach InfoDisen in Wang et al., 2024a , and the right panels report results of the new approach IndiSeek we propose in this paper. Tuning parameters [^2] See Appendix A.2 for an ablation study on performances with different tuning parameter values. are set at $0.01$ for InfoDisen and IndiSeek and at $1$ by default for FactorizedCL.
As we have reasoned when introducing the settings, the ideal modality-specific representations in Setting 1 should depend on and only on the last four coordinates, while in Setting 2, on and only the last two coordinates. The left and the middle panels of both figures suggest that neither FactorizedCL nor InfoDisen learns an ideal representation in either setting, as irrelevant coordinates play non-trivial roles in determining the learned modality-specific representations. Therefore, both SOTA approaches have their limitations in achieving the desired goal of disentangled representation learning. In contrast, the right panels of Figures 1 and 2 show that the new method IndiSeekwe are to propose achieves it in both settings: the learned modality-specific features depend on and only on all coordinates that are independent of those involved in shared features.
1.2 Paper organization
In view of the foregoing limitations of SOTA approaches, in this paper, we introduce a new computationally efficient method IndiSeek, for extracting both shared and modality-specific features, and with a focus on the latter. Based on a reconstruction loss serving as an upper bound for mutual information, IndiSeek is capable of disentangling shared and modality-specific features, while avoiding information loss in each modality at the same time. The effectiveness of IndiSeek is demonstrated on both simulated and real-world datasets, including a single-cell multi-omics dataset CITE-seq, as well as multi-modal benchmark datasets.
The rest of this paper is organized as follows. Section 2 reviews the properties an ideal “shared + modality-specific” feature representation of multi-modal data should have. Section 3 introduces IndiSeek in detail as a practical implementation for seeking such an ideal representation and compares it with SOTA methods in terms of different choices made when approximating information-theoretic quantities. Experiments on a CITE-seq dataset and on MultiBench datasets are presented in Sections 4 and 5, respectively. Section 6 discusses other related works and potential extensions of the proposed method. Technical details and additional numerical studies are deferred to Appendices.
2 Ideal Decomposition of multi-modal data
We consider a fully unsupervised setting with observations from two modalities $X_{1}\in\mathbb{R}^{d_{1}}$ and $X_{2}\in\mathbb{R}^{d_{2}}$. Our goal is to decompose each modality into a shared information component that captures cross-modal information and a modality-specific information component that captures unique variation:
$X_{1}\;\mapsto\;(\underbrace{C_{1}}_{\text{shared}},\underbrace{Z_{1}}_{\text{specific}}),\qquad X_{2}\;\mapsto\;(\underbrace{C_{2}}_{\text{shared}},\underbrace{Z_{2}}_{\text{specific}}).$ (2)
Here $C_{1}$ and $C_{2}$ are $d_{c}$-dimensional [^3] In theory, the dimensions of $C_{1}$ and $C_{2}$ need not match. In practice, these shared latent factors are often obtained from mapping both modalities into some co-embedding space. Therefore, without loss of generality, we assume that $C_{1}$ and $C_{2}$ have the same dimension. shared information components (though not necessarily identical) while $Z_{1}\in\mathbb{R}^{p_{1}}$ and $Z_{2}\in\mathbb{R}^{p_{2}}$ are modality-specific information components.
We formalize the ideal representation / decomposition by requiring the following desiderata:
  1. 1.
    Minimal sufficiency: the shared features preserve all information common to both modalities, i.e.,
    $I(X_{1};X_{2})=I(C_{1};C_{2}).$
    Intuitively, $C_{1}$ and $C_{2}$ should be sufficient for cross-modal prediction tasks. In addition, the shared features should not encode redundant details,
    $I(C_{1};X_{1})=\min\{I(f(X_{1});X_{1}):I(X_{1};X_{2})=I(f(X_{1});X_{2})\},$
    and an analogous identity holds for $C_{2}$. This prevents “over-capturing” modality-specific signals inside the shared space.
  2. 2.
    Independence: modality-specific and shared features should be disentangled,
    $I(Z_{1};C_{1})=I(Z_{2};C_{2})=0.$
  3. 3.
    Complementary information capture: the pair $(C_{i},Z_{i})$ should retain enough information to recover the $i$th modality:
    $I(C_{i},Z_{i};X_{i})\ \text{is maximized for }i=1,2.$
    This ensures that information not explained by shared features is fully captured by the modality-specific components.
Together, these criteria define what it means for a decomposition to be both sufficient and disentangled.
In practice, the extraction of shared features is often carried out with contrastive learning methods such as CLIP, which we also adopt in our framework. Therefore, in the remainder of this work, we focus primarily on the more challenging problem of learning modality-specific representations, assuming that the shared representations have already been obtained.
3 Information-guided disentangled representation seeking (IndiSeek)
Given learned shared features, the desiderata in Section 2 motivate the search for disentangled modality-specific features, but leave open the question of how to realize them. Prior work has proposed an information-theoretic objective (e.g., Wang et al., 2024a, ):
$\max_{Z_{i}}\;I(Z_{i},C_{i};X_{i})\qquad\text{subject to}\quad I(Z_{i};C_{i})\approx 0.$ (3)
This objective reflects the prevailing philosophy: modality-specific features should be maximally informative about their own modality, conditional on the shared component, while being independent of it.
Our position is not to redefine this conceptual objective, but to propose an effective way to implement it in practice. In the following subsections, we describe a two-stage strategy to this end: (1) learning shared features via CLIP, and (2) extracting modality-specific features with a reconstruction-guided objective conditional on learned shared features in step (1). See Figure 3 for an overview of IndiSeek.
3.1 Step 1: Extracting Shared Features via CLIP
As noted at the end of Section 2, shared feature extraction is commonly handled by contrastive learning methods such as CLIP (Radford et al.,, 2021), which we also adopt. Concretely, encoders $(f_{1},f_{2})$ are trained with the InfoNCE loss (Oord et al.,, 2018), which provides a lower bound on the mutual information between $X_{1}$ and $X_{2}$. This encourages $C_{1}=f_{1}(X_{1})$ and $C_{2}=f_{2}(X_{2})$ to capture all cross-modal dependence while avoiding modality-specific redundancy.
Input $\mathbf{X}_{1}$Reconstructed $\widetilde{\mathbf{X}}_{1}$shared $\mathbf{C}_{1}$specific $\mathbf{Z}_{1}$$f_{1}$$h_{1}$$g_{1}$Input $\mathbf{X}_{2}$Reconstructed $\widetilde{\mathbf{X}}_{2}$shared $\mathbf{C}_{2}$specific $\mathbf{Z}_{2}$$f_{2}$$h_{2}$$g_{2}$
Figure 3: IndiSeek: Information-guided Disentangled Representation Seeking.
3.2 Step 2: Extracting Modality-Specific Features
Given the shared features $(C_{1},C_{2})$, we aim to extract the modality-specific components $(Z_{1},Z_{2})$ by operationalizing the constrained problem in Eq. (3), which can be equivalently expressed in Lagrangian form as
$\min_{Z_{i}}\;\underbrace{I(Z_{i};C_{i})}_{\text{disentanglement-enforcing term}}-\quad\lambda\cdot\underbrace{I(Z_{i},C_{i};X_{i})}_{\text{complementary information capture term}}.$ (4)
Here, the minimization over $Z_{i}$ is in fact minimizing over functions $h_{i}(\cdot)$ on $X_{i}$. This form highlights the tradeoff: reducing the dependence of $Z_{i}$ on the shared representation while maximizing the collective informativeness of $(Z_{i},C_{i})$ about $X_{i}$. However, both terms involve mutual information, which is intractable to optimize directly.
To make this objective implementable, we replace each term with a bound that aligns with its optimization direction:
  • For the disentanglement enforcing term $I(Z_{i};C_{i})$, which we aim to minimize, we replace it by an upper bound given by the NCE-CLUB loss (Cheng et al.,, 2020; Liang et al.,, 2023):
    ${\mathcal{L}}_{\rm NCE-CLUB}(Z_{1};C_{1})=\mathbb{E}_{(Z_{1},C_{1})}\left[\log p(Z_{1}\mid C_{1})\right]-\mathbb{E}_{Z_{1},\widetilde{C}_{1}}\left[\log p(Z_{1}\mid\widetilde{C}_{1})\right],$ (5)
    where $p(z_{1}\mid c_{1})$ is the underlying conditional density of $Z_{1}\mid C_{1}$ and $\widetilde{C}_{1}$ is an independent copy of $C_{1}$ from the same marginal distribution. [^4] In implementation, following Liang et al., (2023), we replace conditional distributions by their estimates.
  • For the complementary information capture term $I(Z_{i},C_{i};X_{i})$, which we aim to maximize, we replace it by a lower bound realized through a reconstruction loss that serves as a surrogate for conditional mutual information.
This yields the practical IndiSeek objective that we aim to minimize:
$\mathcal{L}_{\text{{IndiSeek}}}(Z_{i};C_{i},\lambda)\;=\;\mathcal{L}_{\text{NCE-CLUB}}(Z_{i};C_{i})\;+\;\frac{\lambda}{2}\min_{g_{i}}\mathbb{E}\|g_{i}(Z_{i},C_{i})-X_{i}\|^{2}.$ (6)
This formulation encourages $Z_{i}$ to be independent of the shared features while still retaining sufficient information measured via the optimal reconstruction error. Because the objectives are separable across modalities, the extraction procedure can be parallelized.
Justification via reconstruction as MI bound.
When $C_{1}$ is fixed, maximizing ${I(Z_{1},C_{1};X_{1})}=I(Z_{1};X_{1}\mid C_{1}){+I(X_{1};C_{1})}$ is equivalent to maximizing $I(Z_{1};X_{1}\mid C_{1})=H(X_{1}\mid C_{1})-H(X_{1}\mid Z_{1},C_{1})$, which is further equivalent to minimizing the conditional entropy $H(X_{1}\mid Z_{1},C_{1})$. By Fano’s inequality (Cover,, 1999, Theorem 8.6.6), the reconstruction error upper bounds this conditional entropy for any $g_{1}$:
$H(X_{1}\mid Z_{1},C_{1})\;\leq\;\tfrac{1}{2}\log\!\Big(2\pi e\cdot\mathbb{E}\|X_{1}-g_{1}(Z_{1},C_{1})\|^{2}\Big),$
where equality holds if and only if the conditional distribution of $X_{1}\mid(Z_{1},C_{1})$ is Gaussian. This shows why quadratic reconstruction loss is a principled surrogate for conditional mutual information.
3.3 Comparison with two SOTA Methods
Recent approaches to disentangled multi-modal representation learning differ mainly in how they approximate the two terms in Eq. (4): the disentanglement enforcing term $I(Z_{i},C_{i})$, and the complementary information capture term $I(Z_{i},C_{i};X_{i})$.
Factorized Contrastive Learning (Factorized CL, Liang et al.,, 2023).
Like IndiSeek, this method uses the NCE-CLUB loss to upper bound the entanglement enforcing term. However, it replaces the complementary information capture term with an InfoNCE lower bound. For comparison, we consider its task-agnostic, two-step variant: [^5] Liang et al., (2021) also proposes a self-supervised variant of Factorized CL by leveraging task-relevant augmentations, more discussions on which are deferred to Section 6.
${\mathcal{L}}_{\rm FactorizedCL}(Z_{1};C_{1},\tau,\lambda)={\mathcal{L}}_{\rm NCE-CLUB}(Z_{1};C_{1})+\frac{\lambda}{2}{\mathcal{L}}_{\rm infoNCE}(h(C_{1},Z_{1}),X_{1},\tau).$ (7)
Disentangled Self-Supervised Learning (InfoDisen Wang et al., 2024a, ).
This method also relies on InfoNCE for the complementary information capture term, but replaces NCE-CLUB with an orthogonal loss based on von Mises–Fisher assumptions:
${\mathcal{L}}_{\rm InfoDisen}(Z_{1};C_{1},\tau,\lambda)=\mathbb{E}_{(Z_{1},C_{1})}\left[\langle\mu(Z_{1}),\mu(C_{1})\rangle\right]+\frac{\lambda}{2}{\mathcal{L}}_{\rm infoNCE}(h(C_{1},Z_{1}),X_{1},\tau).$ (8)
Both methods inherit limitations: InfoNCE is sensitive to temperature tuning $\tau$ and may underperform as a surrogate for capturing complementary information, while the orthogonal loss assumes linearity and fails to capture nonlinear dependencies. IndiSeek avoids these issues by pairing NCE-CLUB with a reconstruction-based mutual information bound.
4 Experiments with a CITE-seq dataset
In this section, we evaluate the performance of IndiSeek on applications in single-cell biology. We consider the bone marrow CITE-seq dataset from Stuart et al., (2019) which consists of measurements in two modalities on $30672$ individual cells: transcriptome (RNA) and $25$ cell-surface proteins (ADT). We randomly split the dataset into a training set of size $15000$ and a test set of size $15672$. To leverage both RNA and ADT data for annotating cell types, Hao et al., (2021) proposed to cluster cells according to a weighted-nearest-neighbor (WNN) graph with cells as nodes. In its construction, the similarity between a cell $i$ and every other cell $j$ is measured by a weighted average of their respective similarities in the two modalities. The weights used by cell $i$ are in turn determined by the cross-modal predictive powers of its individual modalities after local smoothing. Thus, the RNA weight of each cell can be viewed as a quantification of the importance RNA plays in determining its cell state. The larger, the more important. Based on clustering nodes of the WNN graph, Hao et al., (2021) annotated cells at two granularity levels with $5$ (level-1) and $27$ (level-2) cell types, respectively.
Figure 4: Performance of IndiSeek  in CITE-seq dataset ($\lambda=1$).
In this experiment, we first apply IndiSeek to train neural nets [^6] In this example, representation maps are trained within the class of $5$-layer ReLU neural networks with middle layers of width $50$. that map RNA and ADT data to their disentangled representations, $(C_{\rm RNA},C_{\rm ADT})$ and $(Z_{\rm RNA}$, $Z_{\rm ADT})$, using the training set. We then apply the trained networks to infer disentangled representations of cells in the test set. The ideal disentangled representations should reflect distinctive levels of informativeness that individual modalities have in determining each cell’s annotation. To this end, for each cell $c$, we find its 10 nearest neighbors measured by Euclidean distance in $Z_{\rm RNA}$ and $Z_{\rm ADT}$, respectively, and compare the proportions of neighbors with the same level-2 annotations as $i$ in the two modalities, denoted by $\beta_{\rm RNA}(c)$ and $\beta_{\rm ADT}(c)$, respectively. If the two proportions differ sizably, the modality-specific information in the higher-proportion modality plays a more important role in determining the cell’s annotation, which motivates us to summarize the comparison with a score $\theta(c)=\log(\beta_{\rm RNA}(c)/\beta_{\rm ADT}(c))$ that is monotone increasing with respect to the importance of RNA-specific information.
Figure 5: Comparison of rank correlation metrics across three methods on the CITE-seq dataset.
In the left panel of Figure 4, $27$ level-2 cell types are ranked according to average $\theta(c)$ scores of cells in the test set with respective cell type annotations, in which a higher rank indicates a higher impact of the RNA-specific information in determining the cell type. Each bar is colored according to the coarser level-1 cell type that each level-2 cell type belongs to for easier visual inspection. Different cell types within the T cell family have negative scores, indicating RNA-specific information is less informative compared to ADT-specific counterpart for differentiating different T cell subpopulations, which is in line with the observation that clustering transcriptomics data typically does not delineate T cell subpopulations (Szabo et al.,, 2019; Zheng et al.,, 2021) and gating based on surface protein markers is often needed (Kotliar et al.,, 2025). In contrast, RNA-specific information plays a more important role in distinguishing different progenitor cells than ADT-specific information. To further demonstrate the quality of learned modality-specific representations, we plot the ranks of average $\theta(c)$ values against the benchmark ranks of average RNA weights in Hao et al., (2021) of the 27 level-2 cell types in the right panel in Figure 4, which exhibits a strong correlation that is confirmed by Spearman’s rank correlation (Spearman,, 1961) of $0.922$. As a comparison, we repeated the foregoing experiments across $10$ random seeds while replacing IndiSeek with InfoDisen or FactorizedCL when training the same neural net architecture and the resulting alignment metrics with respect to the results in Hao et al., (2021) are presented in Figure 5. Both alternative methods yielded inferior alignment with the relative importance of the RNA modality in Hao et al., (2021) compared to IndiSeek. To ablate the choice of $\lambda$, in all experiments with real-world data, we report results with varying $\lambda$ on grid values $10^{j}$ for j=3,2,1,0,1,2,3j=-3,-2,-1,0,1,2,3.
5 Experiments with MultiBench datasets
In addition to the application in single-cell multi-omics, we also evaluate the performance of IndiSeek on MultiBench datasets (Liang et al.,, 2021), which include 4 video datasets (MOSI (Zadeh et al.,, 2016), MOSEI (Zadeh et al.,, 2018), UR-FUNNY (Hasan et al.,, 2019), MUStARD (Castro et al.,, 2019)) and Medical Information Mart for Intensive Care III (MIMIC) dataset (Johnson et al.,, 2016). We follow the processing pipeline in Liang et al., (2023) and adopt exactly the same data splitting and feature pre-extraction, in which we focus on the task of binary classification on whether the patient fits any ICD-9 code in group 7 for the MIMIC dataset. We use a smaller Transformer architecture with $2$ heads, $2$ layers, and the intermediate layers with a width $128$. For all methods in comparison, we train the architecture for $2000$ epochs with the same training parameters. We note that our focus here is to benchmark the qualities of representations learned with task-agnostic objective functions. Therefore, for fair comparison along this direction, we do not involve task-related within- or cross-modality data augmentations and hence do not involve self-supervised learning terms in objective functions of all methods in comparison. In contrast, to achieve the optimal performance in a specific task on a particular benchmark dataset, appropriate data augmentation informed by the task and domain knowledge is often helpful.
Following Liang et al., (2023) and Wang et al., 2024a , we focus on the linear probing accuracy for each method with learned representations $(C_{1},Z_{1},C_{2},Z_{2})$ as features, i.e., we concatenate shared and modality-specific features for classification. Here we also compare with the CLIP baseline, where only the shared representations $(C_{1},C_{2})$ are used for linear probing accuracy evaluation. Following the same training procedure as for the CITE-seq dataset, we report the results with three methods in Table 1. For each method, we vary $\lambda$ in $\{10^{j}:j=-3,-2,-1,0,1,2,3\}$ and for each $\lambda$, we run each method with $10$ random seeds. Table 1 presents the averaged accuracy for the best choice of $\lambda$. More details on the performance with varying $\lambda$’s can be found in Appendix A.4.1.
Table 1: Comparison of accuracy on MultiBench datasets (max over $\lambda$, averaged over 10 seeds).
Method MOSI MOSEI UR-FUNNY MUStARD MIMIC
Our method (IndiSeek) 56.81% 73.87% 60.43% 53.91% 65.80%
InfoDisen 55.74% 73.17% 57.67% 53.41% 65.58%
FactorizedCL 55.92% 73.12% 57.61% 53.70% 65.69%
CLIP (baseline) 56.11% 72.49% 56.81% 53.77% 64.60%
From Table 1, we can see that IndiSeek outperms baseline and two SOTA methods across all datasets, and the performances of three disentangled learning methods are comparable in datasets MUStARD and MIMIC, among which, MUStARD ($690$ samples) has a small scale. Moreover, all three disentangled learning methods outperform the CLIP baseline in most of the datasets, which demonstrates the necessity of involving modality-specific features in downstream tasks. In video datasets, such as MOSI ($2199$ samples), MOSEI ($22777$ samples), and UR-FUNNY ($16514$ samples), IndiSeek exhibits more pronounced performance gain in accuracy under the same training configurations, indicating its efficiency in enforcing disentanglement and preserving sufficiency in representations.
To further evaluate the learned task-agnostic representations, we also investigate the multi-task performance of each method on the MIMIC dataset. We consider the tasks of predicting ICD-9 codes for multiple groups in the MIMIC dataset in MultiBench (Liang et al.,, 2021). We compare the performance of all three disentangled learning methods together with CLIP baseline in Table 2, in which we calculate the average accuracy across tasks of predicting the last three ICD-9 codes, and present results in a subset of ICD-9 codes. Complete results for other tasks are deferred to the appendix.
Table 2: Multi-task performance on MIMIC dataset (ICD-9 groups 17–19).
Method group 17 group 18 group 19 Average (3 tasks)
IndiSeek 62.26% 60.92% 69.56% 64.25%
FactorizedCL 61.88% 60.53% 69.43% 63.95%
InfoDisen 61.90% 60.57% 69.50% 63.99%
CLIP 61.33% 60.80% 69.33% 63.82%
6 Extensions and discussion
Guided by information-theoretic principles, we proposed IndiSeek, a method that extracts shared and modality-specific features that are simultaneously sufficient and disentangled. While our focus has been on task-agnostic disentangled representation learning, IndiSeek can be readily extended to task-related settings.
Task-related extensions.
  • With task-specific augmentations. When a downstream task is known in advance, domain knowledge can be used to design within-modality data augmentations. If the augmentations are ideal in the sense that the task-invariant orbit of each observation is fully covered (Liang et al.,, 2023), one may first perform within-modality contrastive learning. The resulting representations retain only task-relevant information. IndiSeek can then be applied to further decompose these task-relevant features into shared and modality-specific components.
  • With side information $Y$. When auxiliary task-related information $Y$ is available, IndiSeek can treat $Y$ as an additional modality and be applied to $(X_{i},Y)$ pairs. The modality-specific features of each $X_{i}$ relative to $Y$ can then guide the design of effective augmentations, after which the procedure above can be repeated on the task-related representations. In the extreme case where $Y$ is the task label itself, each $X_{i}$ can first be reduced to its shared features with $Y$, denoted $\widetilde{X}_{i}$, since $\widetilde{X}_{i}$ contains all information in $X_{i}$ about $Y$. IndiSeek may then be applied to $\widetilde{X}_{1},\widetilde{X}_{2}$ for further decomposition.
Tuning parameter selection.
Our paper also leaves an open question regarding how to select the Lagrange multiplier $\lambda$ in (6) in a principled manner. As the CLUB upper bound and the squared reconstruction loss are at different scales, the optimal $\lambda$ may itself be scale-dependent and vary across datasets. Identifying the optimal monotone transformation of the squared reconstruction loss remains an interesting future direction to enhance empirical performance.
Additional related work.
In the literature, there are other alternatives that learn both shared and modality-specific features in an information-theoretic manner. For example, Dufumier et al., (2024) adopts infoNCE as a proxy for mutual information. However, the framework in Dufumier et al., (2024) cannot ensure disentanglement between shared and modality-specific features, leading to potential redundancy in learned representations. In a concurrent work, Shi et al., (2025) adopts variational bounds for mutual information and entropy regularization as objectives to learn disentangled features. However, similar to Liang et al., (2023) and Dufumier et al., (2024), the objectives are motivated by some task labels $Y$ and hence differ from our goal of learning task-agnostic features.
Acknowledgements
C.M. was partially supported by the National Science Foundation via grant DMS-2311127. Z.M. was partially supported by the National Institutes of Health via grant U01CA294514 and the National Science Foundation via grants DMS-2345215 and DMS-2245575.
Appendix A Additional experimental results
A.1 Details on metrics
Masking-based feature importance.
In simulated experiments, to determine which coordinates are extracted as shared or modality-specific features in $X_{1}$, we consider model-free feature importance based on feature masking related to Robnik-Šikonja and Kononenko, (2008); Zeiler and Fergus, (2014); Li et al., (2016). Concretely, given a black-box architecture $\psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{p}$, to understand which coordinates in each $x\in\mathbb{R}^{d}$ are effective in the output $\psi(x)$, we use leave-one-coordinate-out $x_{-j}$ as inputs for $j\in[d]$ and quantify the difference between and original output and output with masked input $\|\psi(x)-\psi(x_{-j})\|^{2}$. In implementation, for each coordinate $j\in[d]$, we draw $M=1000$ random vectors $\{x_{i}\}_{i\in[M]}$ from uniform distributions on $10,10]^{d}$ and calculate the average
$\zeta_{\psi}(j)=\frac{1}{M}\sum_{i\in[M]}\|\psi(x_{i})-\psi(x_{i,-j})\|^{2}.$ (9)
In this paper, we normalize the vector $\boldsymbol{\zeta}_{\psi}=(\zeta_{\psi}(j))_{j\in[p]}$ to the simplex with unit $\ell_{1}$-norm and use this normalized score $\widehat{\zeta}_{\psi}(j)$ to measure the importance of the $j$th coordinate in $\psi$.
Rank correlation.
Recall that we obtain disentangled representations for RNA and ADT data, $(C_{\rm RNA},C_{\rm ADT})$ and $(Z_{\rm RNA}$, $Z_{\rm ADT})$, using the training set. To this end, for each cell $c$, we find its 10 nearest neighbors measured by Euclidean distance in $Z_{\rm RNA}$ and $Z_{\rm ADT}$, respectively, and compare the proportions of neighbors with the same level-2 annotations as $i$ in the two modalities, denoted by $\beta_{\rm RNA}(c)$ and $\beta_{\rm ADT}(c)$, respectively. We summarize the comparison with a score $\theta(c)=\log(\beta_{\rm RNA}(c)/\beta_{\rm ADT}(c))$ that is monotone increasing with respect to the importance of RNA-specific information, and rank $27$ cell types based on this metric. We use the RNA weights based on weighted k-NN obtained in Hao et al., (2021) as the benchmark and compare our ranks $(R_{i})_{i\in[L]}$ for cell types with the RNA-weights-based ranks $(R^{*}_{i})_{i\in[L]}$ with $L=27$. We adopt Spearman’s $\rho$ correlation as the metric, where for any two vectors of ranks as permutations of $[L]$,
$\rho(\mathbf{R},\mathbf{R}^{*})=\frac{\sum_{i\in[L]}\left(R_{i}-\frac{L(L+1)}{2}\right)\left(R^{*}_{i}-\frac{L(L+1)}{2}\right)}{\sqrt{\sum_{i\in[L]}\left(R_{i}-\frac{L(L+1)}{2}\right)^{2}}\sqrt{\sum_{i\in[L]}\left(R^{*}_{i}-\frac{L(L+1)}{2}\right)^{2}}}.$ (10)
Moreover, when there are no ties in both $\mathbf{R}$ and $\mathbf{R}^{*}$, it holds that
$\rho(\mathbf{R},\mathbf{R}^{*})=1-\frac{6\sum_{i\in[L]}(R_{i}-R^{*}_{i})^{2}}{L(L^{2}-1)}.$ (11)
A.2 Numerical simulations
In this section, we present additional results for the same setting in the main paper with varying $\lambda$’s. Recall the setting with two modalities with $d_{1}=6$ and $d_{c}=2$:
  • Setting 1: Let the observed data be iid copies of $X_{1}{\sim}{\mathcal{N}}(0,I_{d_{1}})$. For each $x\in\mathbb{R}^{d_{1}}$, define the shared representation map
    $f_{1}(x)=0.5A_{f}x+0.2\sin(A_{f}x)+0.2(A_{f}x)^{3}\quad\text{with}\quad A_{f}=(\mathbf{I}_{d_{c}},\mathbf{O})\in\mathbb{R}^{d_{c}\times d_{1}}.$ (12)
    Here, the sine and cubic functions are applied entrywise. The ideal modality-specific features are the last four coordinates of $X_{1}$, which contain all remaining information while being independent of $C_{1}$.
  • Setting 2: Let the observed data be iid copies of $X_{1}$ where $(X_{1})_{\{1,2,5,6\}}\sim{\mathcal{N}}(0,\mathbf{I}_{4})$, $(X_{1})_{3}=0.2\times((X_{1})_{1}+(X_{1})_{2})$, and $(X_{1})_{4}=(X_{1})_{1}\times(X_{1})_{2}$. Thus, the third and fourth coordinates are deterministic functions of the first two, while the others are independent. Let $C_{1}=(X_{1})_{1:d_{c}}$. Here, the ideal modality-specific features are the last two coordinates.
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 6: IndiSeek with varying $\lambda$: feature importance averaged over $50$ simulations (Setting 1).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 7: Factorized CL with varying $\lambda$: feature importance averaged over $50$ simulations (Setting 1).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 8: InfoDisen with varying $\lambda$: feature importance averaged over $50$ simulations (Setting 1).
For Setting 1, given shared features are generated by sine and cubic functions, we can see from Figure 6 and 7 that both Factorized CL and IndiSeek are capable of extracting modality-specific features that are independent of shared ones when $\lambda=0.01$, while IndiSeek tend to be more robust to the choice of $\lambda$ and preserves the desired performance when $\lambda=0.1$. However, in contrast, InfoDisen, due to the limitation of orthogonal loss in handling nonlinear dependence, fails to disentangle modal-specific features from shared ones regardless of the value of $\lambda$.
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 9: IndiSeek with varying $\lambda$: feature importance averaged over $50$ simulations (Setting 2).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 10: Factorized CL with varying $\lambda$: feature importance averaged over $50$ simulations (Setting 2).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 11: InfoDisen with varying $\lambda$: feature importance averaged over $50$ simulations (Setting 2).
From Figures 9, 10, and 11, we can see that similar to the results presented in Section 1.1, with $\lambda$ in the range of $\{10^{j}:j=-2,-1,0,1\}$, Factorized CL fails to enforce independence between shared and modality-specific features; InfoDisen tends to miss modality-specific features with small $\lambda$ and obtain dependent features due to the limitations of orthogonal loss in handling nonlinear dependence. In comparison, with $\lambda=0.1$, IndiSeek can exactly capture the remaining two modality-specific features. Moreover, these figures also reveal the tradeoff between the disentanglement and sufficiency, determined by the value of $\lambda$. With $\lambda=0.01$, more weight is put on the disentanglement between shared and modality-specific features; thus, with a strong regularization on dependence, IndiSeek and InfoDisen tend to miss one modality-specific feature. On the other hand, with $\lambda=10$, the loss puts more weight on reconstruction/sufficiency, and as a result, disentangled methods may extract features that are dependent on the shared ones.
A.2.1 Ablation studies beyond Gaussianity
We also consider ablated variants of Settings 1 and 2 in Section 1.1. Concretely, we generate modality-specific input coordinates (i.e., $(X_{1})_{d_{c}+1:d}$ in Setting 1 and $(X_{1})_{d_{c}+3:d}$ in Setting 2) independently from a discrete distribution
$(X_{1})_{j}\sim{\rm Unif}\{-\sqrt{7}/2,-1,-1/2,1/2,1,\sqrt{7}/2\}.$ (13)
Results for variants (Setting 3 as a variant for Setting 1, and Setting 4 as a variant for Setting 2) are presented in Figures 1217. Performances of different approaches are comparable to those under the Gaussian settings.
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 12: IndiSeek with varying $\lambda$: feature importance averaged over $10$ simulations (Setting 3).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 13: Factorized CL with varying $\lambda$: feature importance averaged over $10$ simulations (Setting 3).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 14: InfoDisen with varying $\lambda$: feature importance averaged over $10$ simulations (Setting 3).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 15: IndiSeek with varying $\lambda$: feature importance averaged over $10$ simulations (Setting 4).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 16: Factorized CL with varying $\lambda$: feature importance averaged over $10$ simulations (Setting 4).
(a) $\lambda=0.01$.
(b) $\lambda=0.1$.
(c) $\lambda=1.0$.
(d) $\lambda=10.0$.
Figure 17: InfoDisenwith varying $\lambda$: feature importance averaged over $10$ simulations (Setting 4).
A.3 CITE-seq dataset implementation details
We consider the bone marrow CITE-seq dataset from Stuart et al., (2019) which consists of measurements in two modalities on $30672$ individual cells: transcriptome (RNA) and $25$ cell-surface proteins (ADT). We randomly split the dataset into a training set of size $15000$ and a test set of size $15672$. We follow the pre-processing in Hao et al., (2021), [^7] https://satijalab.org/seurat/articles/weighted_nearest_neighbor_analysis where we conduct scaling and centering based on moments estimated from the training set only.
In the experiments presented in Figure 4 and Table 5, architectures of representation maps $g$ and $h$ are all $5$-layer ReLU networks with width $50$ and output dimension $50$. Throughout experiments with real-world datasets, we adopt a batch size of $128$ and a learning rate of $10^{-4}$ for training, in which we set the number of epochs to $2000$.
A.4 MultiBench datasets
We evaluate the performance of our method on MultiBench datasets Liang et al., (2021), including video datasets (MOSI (Zadeh et al.,, 2016), MOSEI (Zadeh et al.,, 2018), UR-FUNNY (Hasan et al.,, 2019), MUStARD (Castro et al.,, 2019)) with modalities of text, image, and audio, and the MIMIC dataset (Johnson et al.,, 2016). For video datasets, following Liang et al., (2023); Wang et al., 2024a , we focus solely on vision and text modalities, excluding audio data in these examples. More concretely,
  • MOSI Zadeh et al., (2016): $2199$ YouTube clips with vision and text modalities. Task: binary sentiment classification.
  • MOSEI (Zadeh et al.,, 2018): $23000$ monologue videos with vision and text modalities. Task: binary sentiment classification.
  • UR-FUNNY (Hasan et al.,, 2019): TED talk videos with vision and text modalities. Task: binary humor detection.
  • MUStARD (Castro et al.,, 2019): $690$ TV show clips with vision and text modalities. Task: binary sarcasm detection.
  • MIMIC (Johnson et al.,, 2016): over $36000$ ICU patient records with time-series vitals and tabular data. Task: Following Wang et al., 2024a , we aim at predicting ICD-9 group $7$ membership.
We follow the processing pipeline in Liang et al., (2023) and adopt exactly the same data splitting and feature pre-extraction. We use a smaller Transformer architecture with $2$ heads, $2$ layers, and the intermediate layers with a width $128$.
Table 3: MOSEI results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 72.45% 72.45% 72.50% 73.16% 73.33% 73.65% 73.87% 73.87%
FactorizedCL 72.47% 72.54% 72.64% 73.12% 72.82% 72.58% 72.61% 73.12%
InfoDisen 72.94% 73.17% 72.72% 72.75% 72.87% 72.90% 72.99% 73.17%
CLIP 72.49% 72.49% 72.49% 72.49% 72.49% 72.49% 72.49% 72.49%
Table 4: MOSI results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 56.11% 56.31% 55.93% 56.22% 56.59% 56.81% 56.30% 56.81%
FactorizedCL 55.79% 55.69% 55.92% 54.50% 54.60% 55.38% 54.83% 55.92%
InfoDisen 54.61% 55.07% 54.81% 55.05% 55.04% 55.74% 55.74% 55.74%
CLIP 56.11% 56.11% 56.11% 56.11% 56.11% 56.11% 56.11% 56.11%
Table 5: MUStARD results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 52.54% 52.17% 52.97% 53.91% 53.84% 53.41% 52.46% 53.91%
FactorizedCL 53.70% 53.51% 53.12% 52.10% 51.85% 53.04% 52.79% 53.70%
InfoDisen 50.29% 49.71% 49.64% 52.83% 53.41% 53.33% 53.33% 53.41%
CLIP 53.77% 53.77% 53.77% 53.77% 53.77% 53.77% 53.77% 53.77%
Table 6: UR-FUNNY results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 56.81% 56.82% 57.07% 59.79% 59.79% 60.43% 60.20% 60.43%
FactorizedCL 56.87% 57.61% 57.59% 56.75% 57.39% 57.04% 57.17% 57.61%
InfoDisen 56.30% 55.99% 56.10% 56.44% 56.41% 57.67% 56.89% 57.67%
CLIP 56.81% 56.81% 56.81% 56.81% 56.81% 56.81% 56.81% 56.81%
Table 7: MIMIC results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 64.64% 65.20% 65.44% 65.64% 65.68% 65.65% 65.80% 65.80%
FactorizedCL 64.57% 65.15% 65.09% 65.58% 65.69% 65.52% 65.58% 65.69%
InfoDisen 64.65% 64.80% 64.73% 64.75% 65.02% 65.58% 65.38% 65.58%
CLIP 64.60% 64.60% 64.60% 64.60% 64.60% 64.60% 64.60% 64.60%
A.4.1 Experimental results with varying $\lambda$
In Section 5, we present the accuracies averaged over $10$ seeds with the optimal $\lambda$ for each method. In the following tables, we present more detailed results on the accuracy for each method with different choices of $\lambda$: Table 3 for MOSEI, Table 4 for MOSI, Table 5 for MUStARD, Table 6 for UR-FUNNY, and Table 7 for MIMIC.
A.4.2 Multi-task performance on MIMIC dataset
In this section, we evaluate IndiSeek on all $20$ tasks of predicting ICD-9 groups, and results are summarized in the following table.
Table 8: Multi-task performance on MIMIC (averaged over 10 seeds).
Task ID IndiSeek FactorizedCL InfoDisen CLIP
0 77.79% 77.48% 77.21% 77.21%
1 91.32% 91.28% 91.26% 91.13%
2 70.83% 70.64% 70.76% 70.40%
3 66.64% 66.58% 66.76% 66.49%
4 70.98% 70.94% 70.71% 70.62%
5 72.16% 71.94% 72.07% 71.83%
6 85.87% 85.91% 85.89% 85.73%
7 65.83% 65.87% 65.84% 64.78%
8 65.84% 65.83% 65.42% 65.29%
9 73.15% 73.17% 72.77% 72.50%
10 99.60% 99.60% 99.60% 99.60%
11 89.14% 89.14% 89.14% 89.14%
12 80.93% 80.93% 80.93% 80.93%
13 96.69% 96.68% 96.69% 96.68%
14 69.34% 69.48% 69.19% 69.03%
15 90.98% 90.98% 90.98% 90.98%
16 96.84% 96.84% 96.84% 96.84%
17 62.26% 61.88% 61.90% 61.33%
18 60.92% 60.53% 60.57% 60.80%
19 69.56% 69.43% 69.50% 69.33%
Average 77.83% 77.76% 77.70% 77.53%
In Figure 8, IndiSeek leads the performance in most tasks. Moreover, in many real-world scenarios, downstream tasks are unknown in advance, or learned representations are expected to handle hundreds of tasks simultaneously. In this case, the task-relevant optimality of augmentations could be barely met in Liang et al., (2023), which highlights the flexibility and practicality of task-agnostic disentangled learning.